Your mass is 57kg. Your gravitational potential energy on the top of Temple mountain (altitude 3543m) would be 1.13x10^6 J relative to Lake Louise. What is the altitude of Lake Louise?
1
Expert's answer
2018-05-15T09:30:08-0400
Potential Energy
E=mgh
So
h=E/mg=(1.3×〖10〗^6)/(57×9.8)=2023 m
Therefore the altitude of Lake Louise
h_LL=3543-2023=1520 m